### v1

@@ Abstract
Large Language Models (LLMs), such as the GPT-4 and LLaMA families, have demonstrated considerable success across diverse tasks, including multiple-choice questions (MCQs). However, these models exhibit a positional bias, particularly an even worse “anchored bias” in the GPT-2 family, where they consistently favour the first choice ‘A’ in MCQs during inference. This anchored bias challenges the integrity of GPT-2’s decision-making process, as it skews performance based on the position rather than the content of the choices in MCQs. In this study, we utilise the mechanistic interpretability approach to identify the internal modules within GPT-2 models responsible for this bias. We focus on the Multi-Layer Perceptron (MLP) layers and attention heads, using the “logit lens” method to trace and modify the specific value vectors that contribute to the bias. By updating these vectors within MLP and recalibrating attention patterns to neutralise the preference for the first choice ‘A’, we effectively mitigate the anchored bias. Our interventions not only correct the bias but also improve the overall MCQ prediction accuracy for the GPT-2 family across various datasets. This work represents the first comprehensive mechanistic analysis of anchored bias in MCQs within the GPT-2 models, introducing targeted, minimal-intervention strategies that significantly enhance GPT2 model robustness and accuracy in MCQs. Our code is available at https://github.com/ruizheliUOA/Anchored_Bias_GPT2.
@@ 1 Introduction
Despite their success, recent studies have uncovered a notable flaw: LLMs exhibit a positional bias when tasked with MCQs. Specifically, the performance of these models (e.g., LLaMA (Touvron et al., 2023a, ), LLaMA2 (Touvron et al., 2023b, ), GPT-4 (Achiam et al.,, 2023)) varies significantly depending on the position of the correct answer within the given choices (Pezeshkpour and Hruschka,, 2023; Zheng et al., 2024a, ). We further observe that this vulnerability to positional bias is even worse in the GPT-2 family, ranging from GPT2-Small-124M to GPT2-XL-1.5B (Radford et al.,, 2019). Our investigations reveal that GPT-2 models consistently favour the first choice ‘A’, regardless of the actual position in the input prompt where the correct answer choice is placed, which we term as “anchored bias”.
Previous work primarily mitigated positional bias in MCQ by analysing the impact of different prompt structures (Pezeshkpour and Hruschka,, 2023) or by estimating different datasets’ prior bias based on test samples (Zheng et al., 2024a, ). Such approaches often remain superficial, merely altering the prompt presentation, or are hard to generalise when there is not a large test dataset. Remarkably, there has been a lack of investigation into the internal mechanisms of LLMs that contribute to the anchored bias and strategies to mitigate it without the need for prompt engineering or prior estimation.
@@
Inspired from (Geva et al.,, 2021, 2022) where MLPs can be treated as key-value memories, we use a straightforward yet potent method (Dai et al.,, 2022) to update these critical value vectors in the MLP, effectively mitigate the anchored bias. This adjustment not only mitigates the anchored bias but also enhances the overall MCQ prediction accuracy of the GPT-2 family over 70% averaged across various MCQ datasets and all GPT-2 family. Additionally, we propose a novel strategy to recalibrate the attention patterns by swapping the attention weight between the anchored position and the correct answer choice position. This strategy also mitigates the anchored bias to a certain extent, especially for the classification accuracy improvement of the Indirect Obejct Identification (IOI) dataset (Wang et al.,, 2022) over 90% on GPT2-Medium.
@@ 2 Background: Large Language Models and Mechanistic Interpretability
Figure 2: Left: the architecture of LLMs, where the input prompt is encoded via $W_{E}$, and then the processed information via attention and MLP layer is accumulated back to the residual stream $\mathbf{X}_{\text{mid}}^{\ell}$ and $\mathbf{X}^{\ell}$ at layer $\ell$, respectively. Finally, the residual stream at $L$ layer is unembedded as logits and normalised as a probability distribution for next token prediction. Right: logit lens (Nostalgebraist,, 2020) is used to investigate the contribution of attention pattern and MLP module for the next token prediction.

### v2

@@ Abstract
Large Language Models (LLMs), such as the GPT-4 and LLaMA families, have demonstrated considerable success across diverse tasks, including multiple-choice questions (MCQs). However, these models exhibit a positional bias, particularly an even worse “anchored bias” in the GPT-2 family, where they consistently favour the first choice ‘A’ in MCQs during inference. This anchored bias challenges the integrity of GPT-2’s decision-making process, as it skews performance based on the position rather than the content of the choices in MCQs. In this study, we utilise the mechanistic interpretability approach to identify the internal modules within GPT-2 models responsible for this bias. We focus on the Multi-Layer Perceptron (MLP) layers and attention heads, using the “logit lens” method to trace and modify the specific value vectors that contribute to the bias. By updating these vectors within MLP and recalibrating attention patterns to neutralise the preference for the first choice ‘A’, we effectively mitigate the anchored bias. Our interventions not only mitigate the bias but also improve the overall MCQ prediction accuracy for the GPT-2 family across various datasets. This work represents the first comprehensive mechanistic analysis of anchored bias in MCQs within the GPT-2 models, introducing targeted, minimal-intervention strategies that significantly enhance GPT2 model robustness and accuracy in MCQs. Our code is available at https://github.com/ruizheliUOA/Anchored_Bias_GPT2.
@@ 1 Introduction
Despite their success, recent studies have uncovered a notable flaw: LLMs exhibit a positional bias when tasked with MCQs. Specifically, the performance of these models (e.g., LLaMA (Touvron et al., 2023a, ), LLaMA2 (Touvron et al., 2023b, ), GPT-4 (Achiam et al.,, 2023)) varies significantly depending on the position of the correct answer within the given choices (Pezeshkpour and Hruschka,, 2023; Zheng et al., 2024a, ). We further observe that this vulnerability to positional bias is even worse in the GPT-2 family, ranging from GPT2-Small-124M to GPT2-XL-1.5B (Radford et al.,, 2019). Our investigations reveal that GPT-2 models consistently favour the first choice ‘A’, regardless of the actual position in the input MCQ prompt where the correct answer choice is placed, which we term as “anchored bias” in Fig. 1.
Previous work primarily mitigated positional bias in MCQ by analysing the impact of different prompt structures (Pezeshkpour and Hruschka,, 2023) or by estimating different datasets’ prior bias based on test samples (Zheng et al., 2024a, ). Such approaches often remain superficial, merely altering the prompt presentation, or lacking a comprehensive analysis of fundamental reasons. Remarkably, there has been a lack of investigation into the internal mechanisms of LLMs that contribute to the anchored bias and strategies to mitigate it without the need for prompt engineering or prior estimation.
@@
Inspired from (Geva et al.,, 2021, 2022) where MLPs can be treated as key-value memories, we use a straightforward yet potent method (Dai et al.,, 2022) to update these critical value vectors in the MLP, effectively mitigate the anchored bias. This adjustment not only mitigates the anchored bias but also enhances the overall MCQ prediction accuracy over 70% averaged across various MCQ datasets and all GPT-2 family. Additionally, we propose a novel strategy to recalibrate the attention patterns by swapping the attention weight between the anchored position and the correct answer choice position. This strategy also mitigates the anchored bias to a certain extent, especially for the classification accuracy improvement of the Indirect Obejct Identification (IOI) dataset (Wang et al.,, 2022) over 90% on GPT2-Medium. Finally, we trace the full anchored bias circuit of each GPT2 model, which includes all attention heads and MLPs contributing to this bias.
@@ 2 Background: Large Language Models and Mechanistic Interpretability
Figure 2: Left: the MLP and attention modules of LLMs, where the input prompt is encoded via $W_{E}$, and then the processed information via attention and MLP layer is accumulated back to the residual stream $\mathbf{X}^{\ell}$ at layer $\ell$. Finally, the residual stream at $L$ layer is unembedded as logits and normalised as a probability distribution for next token prediction. Right: logit lens (Nostalgebraist,, 2020) is used to investigate the contribution of attention pattern and MLP module for the next token prediction.
@@
As shown in Fig. 2, the architecture has $L$ layers, and each layer consists of attention and MLP modules, which transform token embeddings to residual streams $(\mathbf{x}_{1}^{\ell},\ldots,\mathbf{x}_{T}^{\ell})\in\mathbf{X}^{\ell}$ at layer $\ell$, where $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$. The residual stream at layer $\ell$ is a place where all attention and MLP modules at layer $\ell$ read from and write to (Elhage et al.,, 2021), and it is updated by the following equation for token $i$ at layer $\ell$:
where $\mathbf{a}_{i,j}^{\ell}$ indicates the attention contribution from token $i$ to token $j$, and $\mathbf{a}_{i}^{\ell}=\sum_{j=1}^{T}\mathbf{a}_{i,j}^{\ell}$. $\mathbf{r}_{i,j}^{l,h}$ indicates the weighted average values where token $i$ attend to token $j$ by head $h$ at the layer $\ell$, and $\mathbf{r}_{i}^{\ell,h}=\sum_{j=1}^{T}\mathbf{r}_{i,j}^{\ell,h}$ (See Appendix C for detailed explanations about attention module).
MLP module is normally treated as key-value memories (Geva et al.,, 2021; Elhage et al.,, 2021; Geva et al.,, 2022; Dai et al.,, 2022), where columns of $W_{\text{in}[:,i]}^{\ell}$ and rows of $W_{\text{out}[i,:]}^{\ell}$ are viewed as keys and values in Fig. 2, respectively. Given the input $\mathbf{x}_{i}^{\ell-1}$, the keys of MLP produce a vector of cofficients $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$, and they weights the corresponding values $\mathbf{v}_{i}^{\ell}$ in $W_{\text{out}}^{\ell}$ (See Appendix C for detailed introduction about MLP):
Logit lens.
3 Preliminaries: Zero-shot Learning with MCQs
Zero-shot learning.
Table 1: The distribution of correct choices on each training dataset. IOI, LD, and Greater-than datasets are manual-synthesised and we did not choose or place the correct choice at A. For test datasets, we only select samples whose correct choices are not ‘A’ to avoid overlap between anchored predictions from GPT2 models and the correct choice. ()($\cdot$)( ⋅ ) indicates the number of choices.
Datasets Train Test A (%) B (%) C (%) D (%) E (%)
IOI (2) - 1000 0 100 - - -
LD (3) - 200 0 50 50 - -
Greater (4) - 1000 0 33.33 33.33 $3$
ARC (4) 1.12k 907 20.82 26.18 25.65 25.29 -
CSQA (5) 9.74k 982 19.60 20.25 19.98 20.38 19.79
Datasets and models.
To comprehensively verify and evaluate the anchored bias of GPT2 family, we consider 5 datasets, which include different numbers of choices from 2 to 5. Indirect Object Identification (IOI) Wang et al., (2022) and Greater-than task (Greater) Hanna et al., (2024): These two datasets have been verified that GPT2 family works well (Wang et al.,, 2022; Merullo et al.,, 2024; Hanna et al.,, 2024). However, we found that GPT2 family immediately fails these tasks if the input prompt is formatted as MCQ in Fig. 1, where the incorrect subject of the last clause or incorrect years is placed in the ‘A’ choice and the prediction is always anchored at incorrect choice ‘A’. Logical Deduction of the Big-Bench (LD) [^2] https://github.com/google/BIG-bench/blob/main/bigbench/benchmark_tasks/logical_deduction/three_objects/task.json Srivastava et al., (2023): LD is a subtask which evaluates three-object logical deduction tasks, and it is used to measure whether model can parse information about multiple choices and their mutual relationships. ARC-Challenge (ARC) Clark et al., (2018) and CommensenseQA (CSQA) Talmor et al., (2019) are commonly-used MCQ benchmarks to evaluate LLMs. For each dataset, we split into 90% Infer. set for anchored bias discovering and mitigation, and 10% Eva. set to evaluate the modified GPT2 model performance. For models, we comprehensively evaluate the GPT2 family, i.e., GPT2-Small-124M, GPT2-Medium-355M, GPT2-Large-774M, and GPT2-XL-1.5B Radford et al., (2019) (See Appendix D for detailed introduction of each dataset).
Evaluation metrics.
4 Discovering Anchored Bias in MCQs
Frequency of anchored bias in GPT2 family across all datasets.
Dist. (%) GPT2-Small GPT2-Medium GPT2-Large GPT2-XL
IOI (2) 45.5 97.4 100.0 85.8
LD (3) 63.0 94.0 100.0 17.0
Greater (4) 32.1 95.0 99.5 98.0
ARC (4) 54.6 91.6 97.6 69.9
CSQA (5) 34.8 81.5 99.6 97.7
As introduced in § 3, we use 5 different datasets with choices from 2 to 5 to investigate the anchored bias. Table 1 shows that the correct choices distribution of ARC and CSQA training dataset is balanced from ‘A’ to ‘E’. In addition, IOI, LD and Greater test datasets are manually synthesised, and we did not choose or place the correct choice at ‘A’. For all test datasets, we only select samples whose correct choice is not ‘A’ to avoid introducing extra bias, i.e., the mix-up between correct prediction and anchored bias of GPT2 family. Based on the randomly sampled test datasets in Table 1, we further calculate the distribution of anchored bias ‘A’ happened within each test dataset when different GPT2 model is used in Table 2. We can find that GPT2-Large and GPT2-Medium have the most serious anchored bias, and GPT2-XL and GPT2-Small have relatively less serious issues. Based on this situation, we mainly focus on investigating test samples which have anchored bias for each dataset. Table 7 in Appendix E shows the number of test samples for each dataset and GPT2 model, where Infer. is used to localise and mitigate anchored bias and Eva. is used to verify the performance of mitigation.
Locating MLP of GPT2 family for anchored bias.
[…]
We calculate the MLP logit difference between anchored bias token ‘A’ and correct choice token B/C/D/E averaged across all layers and datasets for each GPT2 model using Infer. test samples. As shown in Fig. 3, layer 9 in GPT2-Small, layer 20 in GPT2-Medium, layer 34 in GPT2-Large, and layer 37/38/44 in GPT2-XL are dominant layers [^4] In this work, the layer number and head number start from 0. related to anchored bias. In addition, these layers are much closer to the final layer of GPT2, which agrees with Geva et al., (2022); Gurnee et al., (2023)’s finding that higher layers in GPT2 are relevant to semantic concepts or complicated tasks. We also notice that the last one or two layers in each GPT2 model do not have anchored bias at all, and they contribute more logits to correct choice token ‘B/C/D/E’ to ‘A’. However, as the anchored bias logits are accumulated from previous layers, the final one or two layers cannot totally correct this bias.
In the Table 3, vector column demonstrates the specific value vectors $\mathbf{v}^{\ell,n}$ which are responsible for anchored bias. For each value vector, we also calculate how frequently it is recognised as an anchored-bias vector across all datasets for each GPT2 model. We can find that most identified value vectors have more than 50% chance with anchored bias happening across all datasets and different GPT2 models. To further verify whether these value vectors store anchored knowledge bias, we unembeded each value vector logit and selected the top-10 tokens with the highest probability. As shown in Table 3, we can find that most top-10 tokens within each value vector are relevant to ‘A’, e.g., ␣A, A, ␣a, a, etc. This finding proves that some value vectors in $W_{\text{out}}$ of GPT2 family store knowledge bias after pertaining, and these knowledge biases will become anchored bias when the input prompt is formatted as MCQ. In addition, most unembeded tokens are stopwords, e.g., pronouns, articles, prepositions, etc, which also agrees with the findings that “stopwords/punctuation” are commonly distributed in the value vectors of MLP (Geva et al.,, 2022).
Table 3: Identified anchored-bias value vectors $\mathbf{v}^{\ell,n}$ of $n$-row of $W_{\text{out}}^{\ell}$ at layer $\ell$ for each GPT2 model, where the percentage indicates how frequently the specific $\mathbf{v}^{\ell,n}$ is detected as an anchored-bias vector across all datasets, and represents single space within the token because GPT2 tokeniser encodes same word with or without as different token numbers. For each value vector, we further unembedded the top-10 tokens, and most of them are human-interpretable words, which also verify that pretrained GPT2 family has intrinsic anchored bias within $W_{\text{out}}$ (See Appendix F for more unembedded tokens for each GPT2 model).
Model Vector Top-10 Tokens
[…]
$\mathbf{v}^{38,4191}$ (100%) ␣a, ␣an, a, ␣of, ,, ␣and, ., ␣in, an, ␣the
Figure 4: Attention pattern logit difference between anchored bias token ‘A’ and correct tokens (one of B,C,D,E), i.e., $\text{logit}_{T}^{\ell,h}[\texttt{A}](\mathbf{r}_{T}^{\ell,h})-\text{logit}_{T}^{\ell,h}[\texttt{B/C/D/E}](\mathbf{r}_{T}^{\ell,h})$ which is averaged within GPT-2 family across all layers and all datasets. The deeper the blue blocks are at each layer, the more serious the anchored bias is, and vice versa.
Locating attention heads of GPT2 family for anchored bias.
Table 4: The new top-10 tokens of each updated value vector for each GPT2 model (See Appendix G for more new unembedded tokens for each GPT2 model).
Model Updated Vector New Top-10 Tokens
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) ␣B, B, ␣b, ␣C, ␣D, ␣P, ␣L, ␣R, ␣H, ␣F
$\mathbf{v}^{9,2859}$ (61.8%) ␣B, B, ␣b, ␣C, ␣D, ␣L, ␣P, ␣R, ␣G, ␣F
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) ␣C, ␣c, C, ␣A, ␣B, ␣D, ␣F, ␣P, ␣T, ␣G
$\mathbf{v}^{38,4191}$ (100%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣F, ␣T, ␣L, ␣R
As explained in § 2, attention pattern $\mathbf{r}_{i,j}^{\ell,h}$ indicates the weighted average values where token $i$ attend to token $j$ by head $h$ at the layer $\ell$. We use logit lens to analyse the logit difference of final input prompt token contribution between anchored bias ‘A’ and correct choices ‘B/C/D/E’, i.e., $\text{logit}_{T}^{\ell,h}[\texttt{A}](\mathbf{r}_{T}^{\ell,h})-\text{logit}_{T}^{\ell,h}[\texttt{B/C/D/E}](\mathbf{r}_{T}^{\ell,h})$. As shown in Fig. 4, L8H1 and L10H8 in GPT2-Small, L18H12 and L20H5 in GPT2-Medium, L23H8 and L30H0 in GPT2-Large, L31H9 and L34H14555L34H14 indicates layer 34 and head 14 and both of them start from 0. in GPT2-XL are dominant heads related to anchored bias. Those heads are also distributed closer to the final layer in each GPT2 model. We zoom in on the L8H1 and L10H8 attention pattern of the final input token in GPT2-Small using a sample from IOI dataset. As shown in Fig. 5, the final token ‘:’ attends more weights on the anchored bias token ‘A’ than the correct choice token ‘B’, which agrees with our identified attention head using logit difference in Fig. 4. In addition, the full circuit of anchored bias for each GPT2 model can be built based on the MLP and attention logit difference in Fig. 6.
5 Mitigating Anchored Bias in MCQs
Mitigating anchored bias in MLP.
Figure 6: The full circuit of anchored bias for GPT2-Small model, where each attention head and MLP module are selected when MLP and attention pattern logit difference threshold is larger than 4. The percentage within each module indicates the probability of anchored bias across different datasets for GPT2-Small model when threshold is larger than 4 (See Appendix M for the full circuits of other GPT2 models).
According to findings in § 4, we localise the specific value vector in MLP related to the anchored bias. We further aim to solve the following research question: Can we fix the identified value vector in MLP by updating its values and editing the biased knowledge?
[…]
where $\lambda_{1}=1,\lambda_{2}=8$. After updating the corresponding value vector in MLP, we utilise the updated GPT2 model to predict the next token of the same input MCQ prompts. We comprehensively evaluate each updated GPT2 model with the corresponding modified value vector using Infer. and Eva. across all datasets. As shown in Table 5, most updated value vectors achieve high classification accuracy regarding MCQ tasks, even with multiple near 100% or 100% accuracy in $\mathbf{v}^{9,1853}$ of GPT2-Small, $\mathbf{v}^{34,1541}$ of GPT2-Large, $\mathbf{v}^{44,4967}$ and $\mathbf{v}^{38,4191}$ of GPT2-XL. For those value vectors with around a 60-70% chance of anchored bias happening across all datasets and different GPT2 models (i.e., $\mathbf{v}^{9,2859}$ of GPT2-Small, $\mathbf{v}^{20,3713}$ of GPT2-Medium, $\mathbf{v}^{34,2103}$ of GPT2-Large, and $\mathbf{v}^{37,2966}$ of GPT2-XL in Appendix H), the classification accuracy still 68.09% averaged all datasets and all models. This means that the simple and straightforward method (i.e., Eq. 8) is effective, and we do not need to fine-tune the whole GPT2 model to fix the anchored bias. We further unembedded the updated value vectors in Table 4, and it shows that the anchored bias token ‘A’ is significantly removed and the new top-10 tokens for each new value vector are replaced with the correct choices token ‘B/C/D/E’.
Table 5: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model using $\lambda_{2}=8$ (See Appendix H for the ablation study of $\lambda_{2}$ with different values).
Model Vector IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
$\mathbf{v}^{9,2859}$ (61.8%) 44.6 44.2 100.0 100.0 58.8 92.3 93.7 96.2 60.1 56.4
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) 99.4 98.1 97.1 88.5 30.6 70.2 56.8 60.4 26.9 27.5
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) 100.0 100.0 100.0 100.0 100.0 100.0 96.7 96.7 99.7 100.0
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) 98.2 97.8 100.0 100.0 100.0 100.0 90.7 90.8 94.8 94.2
$\mathbf{v}^{38,4191}$ (100%) 100.0 100.0 100.0 100.0 94.9 100.0 97.4 96.9 96.5 95.2
Mitigating anchored bias in attention heads.
[…]
Table 6: The Top-1 classification accuracy changes after attention pattern recalibration for each GPT2 model across all datasets (See Appendix I for other GPT2 model’s results).
Model Head IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
GPT2-XL L31H9 0.0 0.0 9.68 0.0 0.57 0.0 0.53 0.0 0.12 0.0
L34H14 0.0 0.0 12.90 0.0 0.57 0.0 3.33 1.67 0.46 0.0
6 Discussion
Is Few-shot learning helpful?
Based on the comprehensive zero-shot learning MCQ across all datasets, we have a good understanding of how important the MLP and attention head are regarding the anchored bias in the GPT2 family. The following question might be whether few-shot learning can mitigate this anchored bias without updating specific value vectors in MLP or recalibrating the attention head. We conduct an experiment to evaluate GPT2 family across all datasets using 1-shot and 2-shot learning settings. The initial finding is that the anchored bias could be relatively mitigated and average MCQ classification accuracy across GPT2 family is 46.74%, 44.44%, 38.46% and 23.34% under 1-shot learning and 46.52%, 45.70%, 43.22% and 32.62% under 2-shot learning (See Appendix J). This result indicates that GPT2 family still struggles to predict correct choices, especially for GPT2-XL, which needs more investigation in the future.
Is direct value vector updating in MLP harmful to the general ability of GPT2 for other tasks?
We conduct an experiment to evaluate whether direct value vector updating in MLP is harmful to the general ability of GPT2-Small for the original IOI and Greater-than tasks. The experiment shows that the modified GPT2 family can still achieve the average 85.8%, 91.1%, 81.7% and 78.8% accuracy on the original IOI dataset, and 96.1%, 98.0%, 98.5% and 98.4% on the original Greater-than dataset (See Appendix K). Although direct value vector updating in MLP is harmful to the general ability of GPT2 on the original IOI and Greater-than datasets, this model editing approach does not produce serious damage, which matches the findings from Gu et al., (2024). However, we need to develop a better and minimal-harm model editing algorithm in the future.
Is anchored bias sensitive to specific content of input prompts?
We construct two random MCQ datasets which include random concatenated characters and random vocabulary words (see Table 18 and Table 19 in Appendix L). The result shows that anchored bias still happens across different GPT2 models, especially for GPT2-Large and GPT2-XL. This indicates that anchored bias is insensitive to MCQ input prompts, which also confirms our findings that GPT2 family exhibit the anchored bias with significant regularity across various MCQ datasets.
7 Related Work
Mechanistic Interpretability.
Our work is part of Mechanistic Interpretability research, which seeks to decode the internal mechanisms of neural networks (Bereska and Gavves,, 2024; Ferrando et al.,, 2024). Recent studies have deepened our understanding of how transformers manage factual associations. Geva et al., (2023) explores information flow during inference, while Meng et al., 2022a examines knowledge storage in GPT-like models, highlighting their distinct features and using causal mediation analysis to investigate fact recall. Further, Meng et al., 2022b develops an algorithm to enhance memory recall in transformers. Gupta et al., (2023) and Halawi et al., (2023) introduce methods for model editing and behavior analysis, respectively, to mitigate harmful behaviors and adjust commonsense knowledge. Wang et al., (2022) identifies key circuits in attention heads. Our research extends these studies by reformulating IOI as a binary question-answering task to explore positional bias in LLMs.
Positional Bias of LLMs in MCQ.
Several studies have documented the effects of positional bias on LLM accuracy in MCQs. Pezeshkpour and Hruschka, (2023) found a "sensitivity gap" in models like GPT-4, where positional bias can decrease performance by up to 75% in a zero-shot setting, and they improved accuracy with new calibration strategies. Wang et al., (2023) also noted the impact of option order on GPT-4’s scores, enhancing accuracy through a calibration framework including multiple evidence and balanced position adjustments, along with human involvement. Zheng et al., 2024a addressed "selection bias" where LLMs disproportionately favour certain options, introducing a debiasing method, PriDe, that adjusts predictions during inference. Wang et al., (2024) explored performance changes from reordering answer options, confirming that this impacts understanding. Turpin et al., (2024) and Zheng et al., 2024b explore the effects of positional bias in LLMs, showing how it skews Chain-of-Thought generation and evaluator judgments, and emphasize the need for strategies to detect and mitigate these biases. These findings highlight the need for effective strategies to mitigate positional biases, thus enhancing the reliability of LLMs in assessments.
8 Conclusion
Appendix A Limitations
This work mainly focuses on the mechanistic analysis of GPT2 family with the model size from 124M to 1.5B. It is worth comprehensively investigating whether larger open-source LLMs have similar anchored biases, such as LLaMA-7B-65B, LLaMA2-7B-70B, LLaMA3-8B-71B, etc. In addition, different LLM architectural backbones might have different extents of anchored bias, e.g., Mixture of Experts (MoE) and Mamba with selective state spaces. It is meaningful to compare how different MLPs and attention heads are across different LLMs above and why anchored bias disappears if larger LLMs do not have such an issue. Moreover, the knowledge editing approach by directly updating value vectors from MLPs is not optimal as it will introduce some extent of damage to the general ability of GPT2 models. However, how to develop a better and minimal-harm model editing algorithm is an open question (Gu et al.,, 2024), which is worth exploring in the future.
Appendix B Broader Impacts
Mechanistic interpretability of anchored bias for the GPT2 family under the MCQ setting is worth investigating, as it can help us better understand the inner working mechanism of MLPs and attention heads for autoregressive Transformer-based LLMs. The identified MLPs and attention heads leading to the anchored bias can be used to guide the larger LLMs development for safer, less biased and more trustworthy LLMs. Such a mechanistic analysis approach can be extended to other tasks, such as LLMs mathematical reasoning, dialogue generation, and different training methods, such as chain-of-thoughts (CoTs), reinforcement learning from human feedback (RLHF), direct preference optimization. In addition, an adversarial attack might be used for commercial LLM products when this anchored bias is analysed. This also encourages researchers to develop much safer and robust LLMs.
Appendix C Detailed Explanations of LLMs Architecture